OpenStudy (anonymous):

what is the third term in the expansion of (3x-2)^5 A) 1,080X^2 B)270x^3 C)540x^3 D)1,080x^3 please explain how you expanded this to the third term please and thank you:)

OpenStudy (anonymous):

Use binomial expansion

OpenStudy (anonymous):

can this be typed into a calculator

OpenStudy (anonymous):

In my calculator NO

OpenStudy (anonymous):

U Know General Expression T r + 1 = nC r a^(n-r) b^r

OpenStudy (anonymous):

yeah

OpenStudy (raden):

it means 5C2 (3x)^(5-2) (-2)^2 = 5C2 (3x)^3 (-2)^2 = ....

OpenStudy (anonymous):

thanks i got the answer :D

OpenStudy (raden):

yes, u r right ;)